Your organization is planning a conference with 193 people in attendance that need transportation. There are 8 drivers available and 2 types of vehicles, buses and SUV’s . The bus seats 51 people each. The SUV seats 8 people each. How many buses and SUVs will be needed?
Let b= # of buses
Let s= SUVs
Drivers s+b =8
People 51b + 8b = 193

Answer by vleith(2983)

You have one equation in error.
People 51b + 8b = 193
Should be
People 51b + 8s = 193
Now s+b=8
s = 8-b
51b + 8(8-b) = 193
64 + 43b = 193
43b = 129
b = 3
That makes s = 5
What is interesting in this problem is that you are not told to use all the drivers. So, we could use 4 buses and move them all. That saves 4 drivers to so something else. Assuming the cost for gas for one b is less than 4 SUVs, give the other 4 drivers something better to do.